# Breaking Memory Limits: Gradient Wavelet Transform Enhances LLMs Training

### v1

@@ A.2 Experiment Hyperparameters
For the GLUE fine-tuning tasks, we set the hyperparameter to 2 for both GWT and GaLore (enable GWT and GaLore in all modules). The for GWT is chosen from the interval , GaLore is tuned within the interval . For each GLUE task, we use a batch size of 16 (except for CoLA, where the batch size is 32), train for 30 epochs, and set the maximum sequence length to 512. We enable GWT and GaLore across all network modules for all tasks. We present the experimental results of our tuning process in Table 7 as an ablation study. The specific hyperparameters used for the GLUE fine-tuning tasks are summarized in the same Table. All experiments were conducted using a PyTorch [38] random seed of 1234 to ensure reproducibility.
For a fair comparison, we implement GaLore and GWT with an estimated 1/4 optimizer state memory usage. In prior work [12], GaLore was implemented with a rank of 4 and 8, which corresponds to less than 1/4 of the model rank. To achieve the same memory usage with GWT as in GaLore with rank 8, we found that GWT requires approximately a 7-level wavelet transform. Achieving ultra-low memory requirements with GWT, especially beyond the 1/4 model rank, will be explored in future work.
| Haar-2 |
CoLA |
STS-B |
MRPC |
RTE |
SST2 |
MNLI |
QNLI |
QQP |
@@
A.3 Experiment Enviroments
We adopt the BF-16 format for all the pre-training processes to reduce memory usage. All experiments were conducted on 1 to 8 server nodes, each equipped with four NVIDIA 3090 GPUs and Intel(R) Xeon(R) Gold 6226R CPU @ 2.90GHz. The Python version of 3.10, and the PyTorch [38] version of 2.3.0. We use the Pytorch toolbox for our GWT implementation [39].
A.4 Memory and Throughput Estimates
In this section, we provide an additional explanation of the GPU memory usage results reported in the paper. In practical computations, GPU memory usage is influenced by various factors, such as model size, batch size, optimizer states, PyTorch’s memory fragmentation management strategy, and model training strategies (single-GPU or multi-GPU), among others. Here, we estimate the GPU memory usage for the model weights and Adam optimizer states using bfloat16. Specifically, the memory usage for the weights of LLaMA 60M is 0.12G, while the optimizer occupies 0.23G of GPU memory. Since our GWT method primarily targets optimizing the memory usage of the optimizer, the weight memory usage for LLaMA 60M under the Adam optimizer with GWT remains 0.12G. From the perspective of optimizer states, the memory usage of the 2-level GWT should be equal to that of GaLore, which uses only 1/4 of the hidden size as the rank. In our experiments, we applied GWT or GaLore only to the attention and MLP modules of LLaMA. Therefore, the memory usage of the optimizer state also depends on the number of parameters in these two modules. We present the theoretical memory estimate of weight parameters and optimizer states for different methods and models in Table 8.
| Method | 60M | 130M | 350M | 1B |
| Full-Rank∗ | 0.12G/0.23G | 0.25G/0.51G | 0.68G/1.37G | 2.60G/5.20G |
| GWT-2 | 0.12G/0.12G | 0.25G/0.28G | 0.68G/0.54G | 2.60G/1.78G |
| GaLore | 0.12G/0.12G | 0.25G/0.28G | 0.68G/0.54G | 2.60G/1.78G |
| Low-Rank∗ | 0.08G/0.17G | 0.18G/0.37G | 0.36G/0.72G | 1.19G/2.38G |
| LoRA∗ | 0.20G/0.17G | 0.44G/0.37G | 1.04G/0.72G | 3.79G/2.38G |
| ReLoRA∗ | 0.20G/0.17G | 0.44G/0.37G | 1.04G/0.72G | 3.79G/2.38G |
Appendix B Benchmark Details
@@
C4 benchmark. We adopt Colossal Clean Crawled Corpus (C4) [24] benchmark of the English version for all the pre-training tasks. The C4 English benchmark is a colossal, cleaned version of Common Crawl’s web crawl corpus based on the Common Crawl dataset. Includes 305GB of English-language text and is mainly intended to pre-train language models. The validation complexity is obtained by taking the validation loss and exponentiating it with the base of Euler’s number .
@@
Appendix C Availability
All models and benchmarks used in this paper are publicly available. The LLaMA models [11], a series of open-source large-scale models developed by Meta, can be accessed and downloaded from https://www.llama.com/. The network architecture details for these models are provided in Table 5, enabling independent initialization and training. For the fine-tuning task, we use the RoBERTa-base model [25], released by Facebook AI, which can be downloaded from HuggingFace at https://huggingface.co/FacebookAI/roberta-base. The C4 [24] dataset, used for the pre-training tasks, can be downloaded from https://huggingface.co/datasets/allenai/c4 using the Datasets module or via git clone. The GLUE [26] benchmark can be accessed at https://github.com/nyu-mll/GLUE-baselines or through the tools provided by HuggingFace. Finally, the source code to reproduce our results is available at https://github.com/zqOuO/GWT.

### v3

@@ Appendix B Fine-tuning on GLUE
For the GLUE fine-tuning tasks, we set the hyperparameter to 2 for both GWT and GaLore (enable GWT and GaLore in all modules). The for GWT is chosen from the interval , GaLore is tuned within the interval . For each GLUE task, we use a batch size of 16 (except for CoLA, where the batch size is 32), train for 30 epochs, and set the maximum sequence length to 512. All experiments were conducted using a PyTorch [31] random seed of 1234 to ensure reproducibility.
| CoLA | STS-B | MRPC | RTE | SST2 | MNLI | QNLI | QQP | Avg. | |
| Adam | 63.81 | 91.26 | 92.25 | 79.06 | 94.57 | 87.29 | 92.18 | 92.28 | 86.58 |
| GaLore | 61.32 | 91.13 | 92.41 | 77.25 | 94.03 | 86.77 | 92.56 | 91.77 | 85.90 |
| APOLLO | 61.07 | 90.70 | 92.00 | 78.33 | 93.57 | 87.21 | 92.27 | 91.85 | 85.87 |
| GWT | 62.57 | 91.16 | 93.26 | 79.42 | 94.26 | 87.37 | 92.53 | 91.94 | 86.56 |
| CoLA | STS-B | MRPC | RTE | SST2 | MNLI | QNLI | QQP | |
@@
Appendix C Hyperparameter Study
In addition to the hyperparameters required by the original optimizer, several additional hyperparameters are needed to apply our GWT method. The hyperparameter stands for the scale factor. Specifically, the scale hyperparameter determines the magnitude of the update for the GWT module of ( for the modules without GWT). The hyperparameters depend on the GWT basis used in the process (e.g., Haar, DB2, Rbio2.6, etc.). The choice of filter impacts the transformation applied during the wavelet decomposition and reconstruction. stands for the GWT level, 1 denotes 1-level GWT which can reduce the optimizer memory up to 50%, and 2 denotes 2-level GWT which reduces the memory usage up to 75%, respectively. These levels allow for varying degrees of memory reduction depending on the depth of the wavelet decomposition applied. Larger can bring higher memory reduction but influence the throughput. These additional hyperparameters allow for more flexibility in applying GWT to different optimization problems while controlling the trade-off between memory usage and update precision. In this section, we investigate the effects of the hyperparameters and GWT level .
Scale . We evaluate Adam with 2-level Haar GWT using different values of at a fixed learning rate (). As shown in Figure 6(a) and 6(b), we observe that our GWT is invariant to the hyperparameter. For values larger than , the final experimental results remain relatively stable and do not differ significantly. Based on the network initialization strategy we used, most modules in the network have a learning rate greater than 0.001 (). In contrast to Adam, which often requires a learning rate smaller than 0.001 to avoid loss spikes, GWT shows significantly better robustness to higher learning rates. For fine-tuning tasks, tuning is equivalent to adjusting the learning rate () hyperparameter based on our strategy. A detailed study of the learning rate hyperparameters (,) on fine-tuning tasks is presented in Appendix A.2.
GWT level . We investigate the effects of the GWT level hyperparameter , with the corresponding validation perplexity curves shown in Figure 5. We found that the value of has little impact on the final experimental results, and lower values of are only slightly better than others. One question that arises from the experimental results is whether the approximation coefficients are not crucial when using wavelets to compress gradients in LLM training, but rather the detail coefficients. The main impact of seems to be on memory usage and throughput. The throughput results and optimizer memory estimate for different GWT levels are summarized in Table 12, demonstrating the trade-off between memory efficiency and model performance. Higher wavelet levels improve memory usage, but this may come at the cost of slightly higher validation perplexity and slower training speed.
| Methods | 60M | 130M | 350M | |||
| Memory | Tokens/s | Memory | Tokens/s | Memory | Tokens/s | |
| GWT-1 | 0.16G | 95.8K | 0.36G | 46.0K | 0.81G | 12.7K |
| GWT-2 | 0.12G | 91.9K | 0.28G | 45.1K | 0.54G | 12.1K |
| GWT-3 | 0.10G | 90.1K | 0.24G | 44.8K | 0.40G | 11.5K |
| GWT-4 | 0.09G | 84.2K | 0.22G | 43.8K | 0.33G | 10.8K |
| GWT-5 | 0.08G | 83.5K | 0.21G | 41.5K | 0.29G | 9.25K |
Wavelet filter . In the main text, our GWT method uses the Haar wavelet basis by default. Different wavelet bases capture information in distinct ways, which can result in variations in compression performance. To investigate the impact of different wavelet bases, we evaluated GWT’s pretraining performance and memory overhead using several orthogonal wavelet bases, including Bior (Biorthogonal Wavelets), SYM (Symmetric Wavelets), and Coif (Coiflets). The experimental results, presented in Table 13, show that the choice of wavelet basis has minimal effect on the overall experimental outcomes.
| Wavelet | 60M | 130M |
| Haar-2 | 29.35 (0.24G) | 22.47 (0.52G) |
| DB-2 | 29.60 (0.24G+) | 22.59 (0.52G) |
| Bior-2 | 29.34 (0.24G+) | 22.37 (0.52G+) |
| SYM-2 | 29.42 (0.24G) | 22.45 (0.52G) |
| Coif-2 | 29.24 (0.24G+) | 22.40 (0.52G) |
Appendix D Adam with Module-wise Learning Rate
One key learning rate strategy employed in GWT is the assignment of different learning rates to distinct parameter modules. Specifically, the overall learning rate is partitioned such that the Attention and MLP modules are updated using a scaled learning rate of , while the rest of the model uses the base learning rate . This modular learning rate strategy is commonly adopted in memory-efficient training algorithms, including LoRA [21], GaLore [54], Fira [6], and APOLLO [56]. We incorporate this strategy into GWT as a standard configuration. The method of using different learning rates for different network modules in Adam optimizer has been studied in related works [12, 14]. Here, we provide a supplement as an additional ablation experiment to the original text.
In this section, we first apply the same modular learning rate scheme to the Adam optimizer to examine whether it also benefits from this design, and to evaluate whether GWT can still match or surpass Adam’s performance, as demonstrated in the main text. Following standard practice in other memory-efficient methods, we use a learning rate of for the Attention and MLP modules, while keeping for the remaining modules. To measure the benefit of this configuration, we compare it with Adam using a unified learning rate of 2.5e-3, selected from the range [1e-3, 2.5e-3, 5e-3, 7.5e-3, 1e-2] based on best performance. The learning curves are shown in Figure 7. We then apply a uniform learning rate ( = 1.0) to GWT and evaluate its performance under this setting, with results reported in Figure 8.
Experimental results show that, under our setup, Adam significantly benefits from the modular learning rate strategy, achieving clearly better performance than with a single global learning rate. Even with this enhanced setting, GWT achieves a final validation perplexity (PPL) close to that of Full-Adam. This partly explains why many memory-efficient training methods can outperform full-rank Adam—the modular learning rate strategy plays a key role. It also suggests that Attention and MLP modules may require smaller learning rates than other parts of the model. These findings raise an important question: Do different modules in LLMs inherently demand distinct learning rates? And could explicitly integrating modular learning rate strategies into Adam lead to further gains?
Appendix E Discussion
Appendix F General Optimizer with GWT
This equation captures the integration of wavelet transform-based adjustments into the optimizer’s update rule. To this end, we have developed a unified update schedule for optimizers with GWT based on linear formulation and demonstrated that GWT can be directly applied to existing optimizers with heavy-state memory usage. Therefore, we incorporate our GWT into Adam-mini [53] and MUON [24] optimizers for LLaMA pre-training tasks, and present the experimental results in Figure 4.
Appendix G Benchmark Details
@@
C4 benchmark. We adopt the Colossal Clean Crawled Corpus (C4) [34] benchmark of the English version for all the pre-training tasks. The C4 English benchmark is a colossal, cleaned version of Common Crawl’s web crawl corpus based on the Common Crawl dataset. Includes 305GB of English-language text and is mainly intended to pre-train language models. The validation complexity is obtained by taking the validation loss and exponentiating it with the base of Euler’s number .
MMLU Benchmark. We evaluate the performance of LLaMA3.2-3B [15], Qwen2.5-3B [51], and Gemma3-4B [41] on the Multi-task Language Understanding (MMLU) benchmark [18]. The MMLU benchmark is designed to assess the knowledge acquired by models during pretraining, focusing on their performance in zero-shot and few-shot settings. It spans 57 subjects, including areas in STEM, the humanities, social sciences, and more.